OpenStudy (anonymous):

Polar coordinates of a point are given. Find the rectangular coordinates of the point. (-5, -180°)

OpenStudy (anonymous):

\[x = r\cos\theta\] \[y = r\sin\theta\] So, (x,y) = (5,0)

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} (&-5, &-180^o)\\ &\uparrow &\uparrow \\ &r&\theta \end{array}\qquad y=rsin(\theta)\qquad x=rcos(\theta)\)